OpenStudy (anonymous):

Variable x is 7 more than variable y. Variable x is also 1 less than y. Which of the following pairs of equations best models the relationship between x and y? A.x = 1y x = y + 7 B.x = y + 7 x = y − 1 C.y = x + 8 y = x − 1 D. y = 7x y = x + 1

OpenStudy (anonymous):

@israa88

OpenStudy (anonymous):

do you know how to do it @israa88

OpenStudy (anonymous):

i would go with B

OpenStudy (anonymous):

because

OpenStudy (anonymous):

they say that x=7 + y x=x-1

OpenStudy (anonymous):

B

OpenStudy (anonymous):

does that make sense ? @tykeke878

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

need help on anything else ?

OpenStudy (anonymous):

yes I tagged you in it